a) When is the product of 2 binomials a trinomial? Use examples
b) When is the product of 2 binomials a binomial? Use examples

Answer by stanbon(75887)

a) When is the product of 2 binomials a trinomial? Use examples
If you use FOIL, you get a trinomial if O and I are "like" terms.
(x+2)(x+3) = x^2 + 3x + 2x + 6 = x^2 = 5x + 6
-----------------------------------------------------------------------
b) When is the product of 2 binomials a binomial? Use examples
If you use FOIL, you get a binomial if the sum of the "like" terms is zero
(x+2)(x-2) = x^2 -2x + 2x -4 = x^2 + 0x - 4 = x^2 - 4
===========================================================
Cheers,
Stan H.